# NestedSGX: Bootstrapping Trust to Enclaves within Confidential VMs

### v1

@@ 3.2. Memory isolation
Booting NestedSGX. Upon receiving a request to launch a CVM, the platform proceeds to load the VM image and cryptographically measures its contents. During the initialization process, the entire gPA is divided into two distinct parts. By configuring the RMP attributes, we establish a dedicated region of guest memory exclusively reserved for VMPL0,, while the remaining gPA regions are reserved for VMPL1. The memory for VMPL0 is further divided into two parts, one for the security monitor and one for the enclaves.
@@
Isolation of enclaves. It is also important to prevent an enclave from tampering with the security monitor or the guest OS. This ensures protection against potentially malicious enclave (referred to as enclave malware) or enclaves that may have vulnerabilities that could be exploited. One approach to address this concern is to place the enclave at a distinct VMPL level, separate from VMPL0 or VMPL1, and enforce the appropriate VMPL permissions to contain the enclave code. However, the transition is emulated by the security monitor, necessitating at least two VMPL switches. As a result, transitioning between the application and the enclave becomes costly.
@@
Preventing memory mapping attacks. In Intel SGX, the enclave and application share the same page table, and the hardware prevents memory mapping attacks by performing additional security checks during a page table walk, using the EPCM to ensure that the memory mappings are correct.
Without hardware support, it is necessary to prevent the guest OS from manipulating the mappings of the enclave. NestedSGX employs a shadow page table scheme: the application’s page table is managed by the guest OS, while the enclave’s page table is managed by the security monitor. In situations where the application and enclave need to share memory, such as for parameter passing, NestedSGX maps the shared memory to both page tables with the same gVA-to-gPA mapping. The mapping of the parameter buffer remains fixed throughout the entire lifetime of the enclave. Consequently, we do not need to maintain the synchronization of the two page tables.
3.3. Enclave life cycle management
In NestedSGX, the life cycle of an enclave closely resembles that of SGX. Initially, the application invokes the ECREATE leaf function to create the enclave. Subsequently, EPC pages necessary for the enclave runtime, such as code sections, data sections, and Thread Control Structure (TCS) pages, are added using EADD and measured using EEXTEND leaf functions. Once all the required EPC pages are added, the enclave is initialized with EINIT, and the final measurement is performed. Upon completion of the enclave’s execution, the EPC pages are reclaimed using the EREMOVE leaf function, and the enclave is destroyed. NestedSGX achieves compatibility with Intel SGX by transitioning to the security monitor, where the security monitor emulates the instructions strictly adhering to the SGX specifications. Notably, the SGX control structures are stored in the memory of the security monitor, while the code and data pages of the enclave are stored in the memory specifically reserved for the EPC.
Synchronous enclave entry and exit. Once an enclave is initialized, the App can enter the enclave using the EENTER leaf function and jump to the enclave’s code for execution. Since VMPL switches can only be executed in the privileged mode (Sec. 2.2), NestedSGX provides the NestedSGX-driver within the guest OS, which accepts the request from the App and facilitates the VMPL switch. During this process, the NestedSGX-driver receives the type of SGX leaf functions, while the parameters are saved in the parameter buffer. The NestedSGX-driver then switches the execution to VMPL0 and transfers control to the security monitor.
As shown in Figure  4, during the emulation of EENTER, the security monitor switches to the enclave’s page table and enters the enclave entry point specified by the TCS using the sysret instruction. The enclave can then return to the App using the EEXIT leaf function, which is also emulated by the security monitor. Specifically, the enclave returns to the security monitor using the syscall instruction. Subsequently, the security monitor switches the execution to the NestedSGX-driver running at VMPL1. The NestedSGX-driver restores the App’s page table and returns control to the App using the sysret instruction.
Asynchronous enclave entry and exit. During enclave execution, faults and exceptions may occur. In accordance with the SGX model, these faults and exceptions are supposed to be handled by the guest OS first. In NestedSGX, when a fault or exception occurs within the enclave, it traps to the fault and exception handler within the security monitor (Figure 4). The security monitor then emulates the Asynchronous Exit (AEX) event by saving the context of the enclave in the state saving area (SSA). Subsequently, the security monitor switches the vCPU to the NestedSGX-driver, which operates at VMPL1. It fills the instruction pointer (RIP) to the trampoline area (AEP) specified in the TCS and invokes the handler of the guest OS. After the OS handles the fault or exception, it uses the iret instruction to return to the AEP. The AEP, in turn, resumes the execution of the enclave using the ERESUME leaf function. The emulation of the ERESUME leaf function is similar to that of EENTER, except that the enclave’s execution context is restored from the SSA.
It’s worth noting that if the untrusted NestedSGX-driver chooses not to fill RIP with AEP, this decision does not introduce any new security issues. This is because the control flow outside the enclave is already vulnerable to manipulation within the SGX model and is not inherently trusted.
In NestedSGX, faults and exceptions are firstly trapped to the security monitor, enabling easy detection of suspicious behaviors such as abnormal interrupts caused by page fault-based attacks. Conceptually, NestedSGX can also support the recent AEX-notify feature (Constable et al., 2023), as well as enclave dynamic memory management (EDMM) (McKeen et al., 2016) in SGX, although the implementation of these features is planned for future work.
@@ 3.4. Measurement, attestation and sealing
When the CVM image is loaded and measured, the page attributes, including the access permissions of all VMPLs, are included in the measurement. Any attempts to modify the code and data in VMPL0 and VMPL1, or to manipulate the access permissions of guest pages during CVM initialization, will be reflected in the measurement. After the CVM is initialized, software operating at any VMPL level can initiate an attestation report request by sending a message to the AMD-SP firmware. The request structure comprises the VMPL level and 512 bits of space for user-provided data, which will be incorporated into the attestation report signed by the AMD hardware.
@@
Chain of trust. Specifically, when the CVM is first initialized, the security monitor randomly generates a key pair, which serves as the attestation identity key (AIK) of NestedSGX. To establish the binding between NestedSGX ’s AIK and the CVM, we include the digest of the public part of the attestation key in the user-data field of the attestation report request (Figure 5), while the private key never leaves the memory of the security monitor.
The security monitor initiates the generation of an SEV-SNP attestation report by dispatching a SNP_REPORT_REQ message to the AMD-SP hardware. The request message between the security monitor and AMD SP is encrypted using the appropriate VM Platform Communication Key (VMPCK) for VMPL0. Upon receiving the encrypted message, the AMD-SP hardware decrypts it, verifies its integrity, and subsequently responds with an attestation report, which is signed with the SEV-SNP attestation key (Versioned Chip Endorsement Key, VCEK). The report includes the platform measurement, the security monitor measurement, the VMPL level, and the public key of NestedSGX’s AIK. The NestedSGX AIK is later utilized to sign the report of the enclave. This signed enclave report, along with the SEV-SNP attestation report, collectively constitutes the final attestation report in NestedSGX.
@@ 4. Implementations
Security monitor
Emulation of SGX data structures, instructions and AEX
5,500
NestedSGX-driver
Handling switches between App and security monitor
1,400
Intel SGX SDK
Replacing SGX instructions with IOCTL and system calls
500
Total 7,400
@@
Security monitor. We based our implementation on the main branch of SVSM without current privilege level (CPL) support, and added our own support for running user-space enclaves. The security monitor emulates SGX leaf functions based on requests from the NestedSGX-driver. To accommodate these requests, a new type of SGX protocol was introduced, extending the SVSM framework. The security monitor effectively enters a request loop within the SVSM framework, allowing it to continue execution upon the arrival of subsequent requests. The security monitor handles a memory pool acting as the EPC, and faithfully emulates SGX data structures and operations in strict adherence to the SGX specification. We registered customized fault and exception handlers for the enclave, emulating the AEX event. Furthermore, we configured the system call handler by setting the LSTAR MSR to handle the EEXIT leaf function.
@@
With the aforementioned design, the majority of SGX programs can run on NestedSGX without requiring modifications to the source code. Furthermore, to facilitate the porting of existing applications, we have ported the Rust SGX SDK (Wang et al., 2019) to the NestedSGX platform. The support of popular library OSes such as Occlum (Shen et al., 2020) is under development.
@@ 5. Security Analysis
While the enclave operates at VMPL0, it is prohibited from directly obtaining the SEV-SNP attestation report. Otherwise, the enclave may incorporate a counterfeit AIK into the report, which could compromise the integrity of the trust chain, particularly if the enclave deliberately exposes the private key of the AIK. This protection is enforced by the fact that SNP_REPORT_REQ can only be initiated from the kernel mode, which is the security monitor. More specifically, SNP_REPORT_REQ necessitates shared memory with the host VMM for passing request and response messages, but we intentionally disable shared memory between the enclave and the host VMM.
The hardware also prevents the guest OS (operating at VMPL1) from obtaining an SEV-SNP attestation report for VMPL0. Specifically, the desired VMPL is provided by the guest in the SNP_REPORT_REQ message, and the guest can only generate attestation reports for VMPLs that are greater than or equal to the current VMPL (AMD, 2023d, Sec. 7.3).
@@ 6.1. Experimental setup
We deployed NestedSGX on a server with two AMD EPYC 7543 CPUs (two threads per core, total of 128 logical cores) with 64 GB DDR4 RAM. Every CVM was allocated 4 vCPUs and 4 GB RAM, and ran Ubuntu 22.04 with kernel version 6.5.0-snp-guest (svsm-preview-guest-v3 branch provided by AMD). We configured 512 MB gPA for the security monitor and the enclaves. The modified Linux SGX SDK is based on version 2.21. We used the main branch of the linux-svsm framework (commitID: 8d518f1). The SVSM framework and the security monitor were compiled with Rust in the debug mode (-O0) [^3] Currently, the linux-svsm framework cannot be compiled successfully with optimization levels other than -O0.. All programs were compiled with GCC 11.4.0 and the same optimization level (-O2). All evaluations were performed in the single-threaded mode. The baseline were conducted on the same AMD server running in the guest OS directly (native mode) and in the SGX SDK simulation mode (sim mode), which do not provide any security guarantees. Specifically, we used the same code base under the same optimization level and compiling options.
@@ 6.2. Micro-benchmarks
Table 2. Latencies of SGX primitives. A VMPL switch round-trip costs about 19,400 cycles on our platform, while the SGX SDK routine costs 4000–5500 cycles. The rest is the cost of our unoptimized implementation of the security monitor in emulating SGX instructions (about 12,000 cycles). The Intel platform was equipped with an Intel Xeon E3-1270 v6 CPU and 64 GB RAM.
Intel SGX hardware mode
Simulation (insecure) mode
NestedSGX
ECALL
14,437
5,500
37,829
OCALL
12,586
3,927
35,281
Context switches. We measured the latency of ECALLs and OCALLs in NestedSGX. The test runs empty edge calls with no explicit parameters 100,000 times and takes the median value. As shown in Table 2, the latency of context switches in NestedSGX is about 2×2$\times$2 × higher than that on Intel SGX. Notably, existing works on Intel SGX, including EleOS (Orenbach et al., 2017), HotCalls (Weisse et al., 2017) and Switchless calls (Tian et al., 2018) can be applied to NestedSGX to reduce the cost brought by context switches.
@@
Linux/Unix nbench. Linux/Unix nbench is a benchmark suite that focuses on evaluating the performance of a computer system’s CPU, FPU, and memory system (Mayer, 2017). It specifically excludes I/O operations and system calls from its measurements. We utilized an adapted version of nbench, namely SGX-NBench (sgx, 2017) to evaluate the computation and memory performance of NestedSGX. Notably, the adapted SGX-NBench involves a number of ECalls during the evaluation, making a fair comparison with nbench running on the native CVM difficult. Figure 5(a) demonstrates that NestedSGX introduces an overhead of approximately 2.1% on average over the baseline running in SGX simulation mode, which we believe mostly attributes to the timing difference of context switches.
Lmbench. Lmbench is a benchmark suite used to evaluate different aspects of system behavior, such as memory access speed, system call latency, and disk I/O performance. To measure the I/O performance of NestedSGX, we utilized the SGX version of Lmbench called SGX-bench (Hasan et al., 2020). SGX-bench includes three test cases: bw_mem for memory bandwidth evaluation, and lat_rand and lat_syscall for system call evaluation. We conducted 10,000 repetitions of each test case and calculated the average values. The memory buffer size for bw_mem was set to 2 MB. The results are shown in Figure 5(b), showing that the memory bandwidth is 98.7% of the baseline and the measured system calls have negligible overhead. In evaluating the system call latency, the null system call was observed to consume approximately 0.0853 ms, making the context switches nearly imperceptible. These findings align with the evaluations conducted on Intel SGX (Hasan et al., 2020).

### v2

@@ 3.3. Memory isolation
Figure 4. The guest physical address space is divided into 3 parts: the secure memory used by the security monitor, the EPC memory, and the normal memory.
The entire guest physical address space of the CVM is divided into 3 parts (Figure 4): the secure memory used by the security monitor, the secure memory used by the enclaves (i.e., EPC), and the normal memory used by the guest OS and untrusted applications. The access to the secure memory is restricted solely to VMPL0, whereas the normal memory can be accessed by both VMPL0 and VMPL1. Notably, the SGX control structures are stored in the memory of the security monitor, while the enclave’s code and data pages are stored in the memory reserved for the EPC.
Booting NestedSGX. Upon receiving a request to launch a CVM, the platform proceeds to load the VM image and cryptographically measures its contents. During the initialization process, the entire gPA is divided into two distinct parts. By configuring the RMP attributes, we establish a dedicated region of guest memory exclusively reserved for VMPL0, while the remaining gPA regions are reserved for VMPL1. The memory for VMPL0 is further divided into two parts, one for the security monitor and one for the enclaves.
@@
EPC memory management. In Intel SGX, the enclave and application share the same page table, and the hardware prevents memory mapping attacks by performing additional security checks during a page table walk, using the EPCM to ensure that the memory mappings are correct. Without hardware support, it is necessary to prevent the guest OS from manipulating the mappings of the enclave. NestedSGX employs a shadow page table scheme: the application’s page table is managed by the guest OS, while the enclave’s page table is managed by the security monitor. In situations where the application and enclave need to share memory, such as for parameter passing, NestedSGX maps the shared memory to both page tables with the same gVA-to-gPA mapping. The mapping of the parameter buffer remains fixed throughout the entire lifetime of the enclave. Consequently, we do not need to maintain the synchronization of the two page tables.
In NestedSGX, the security monitor is responsible for managing the EPC memory. It maintains two lists: a free list and a used list, containing all the EPC pages. Similarly, an EPCM data structure is utilized to track the state of each EPC page. During enclave initialization, when an EPC page needs to be added to the enclave (via EADD), the security monitor selects a page from the free list and constructs the corresponding page table entry (PTE).
In cases when the enclave’s memory accesses result in a page fault, the CPU triggers a trap to the security monitor. The security monitor forwards the page fault information, including the address of the faulting page, to the NestedSGX-driver. If the NestedSGX-driver determines that a page frame needs to be allocated for the enclave, it requests the security monitor to allocate a physical page within the EPC. Additionally, the security monitor constructs the corresponding PTE in the enclave’s page table. Once the page fault is successfully handled, the security monitor restores the enclave’s execution context. Currently, NestedSGX does not support swapping the enclave’s pages to regular memory or disk storage.
Isolation of enclaves. It is also important to prevent an enclave from tampering with the security monitor or the guest OS. This ensures protection against potentially malicious enclave (i.e., enclave malware) or enclaves that may have vulnerabilities that could be exploited. One approach to address this concern is to place the enclave at a distinct VMPL level, separate from VMPL0 or VMPL1, and enforce the appropriate VMPL permissions to contain the enclave code. However, the transition is emulated by the security monitor, necessitating at least two VMPL switches. As a result, transitioning between the application and the enclave becomes costly.
@@ 3.4. Measurement, attestation and sealing
When the CVM image is loaded and measured, the page attributes, including the access permissions of all VMPLs, are included in the measurement. Any attempts to modify the code and data in VMPL0 and VMPL1, or to manipulate the access permissions of guest pages during CVM initialization, will be reflected in the measurement. After the CVM is initialized, software operating at any VMPL level can initiate an attestation report request by sending a message to the AMD-SP firmware. The request structure comprises the VMPL level and 512 bits of space for user-provided data, which will be incorporated into the attestation report signed by the AMD hardware. The local/remote attestation and sealing mechanisms of NestedSGX closely mirror those employed by Intel SGX. This is accomplished by emulating the corresponding instructions (i.e., EGETKEY and EREPORT) within the security monitor.
@@
Chain of trust. When the CVM is first initialized, the security monitor randomly generates a key pair, which serves as the attestation identity key (AIK) of NestedSGX. The AIK is only required to be generated once and can be reused across different CVM boot cycles. To establish the binding between NestedSGX’s AIK and the CVM, we include the digest of the public part of the attestation key in the user-data field of the attestation report request (Figure 5), while the private key never leaves the memory of the security monitor.
The security monitor initiates the generation of an SEV-SNP attestation report by dispatching a SNP_REPORT_REQ message to the AMD-SP hardware. The request message between the security monitor and AMD SP is encrypted using the appropriate VM Platform Communication Key (VMPCK) for VMPL0. Upon receiving the encrypted message, the AMD-SP hardware decrypts it, verifies its integrity, and subsequently responds with an attestation report, which is signed with the SEV-SNP attestation key (Versioned Chip Endorsement Key, VCEK). The report includes the platform measurement, the security monitor measurement, the VMPL level, and the public key of NestedSGX’s AIK. The NestedSGX AIK is exclusively accessible to the quoting enclave (QE) and is subsequently used to sign the report of the enclave. This signed enclave report, along with the SEV-SNP attestation report, collectively constitutes the final attestation report in NestedSGX.
@@ 4. Implementations
Security monitor Emulation of SGX data structures, instructions and AEX 5,500
NestedSGX-driver Handling switches between App and security monitor 800
SGX SDK & Occlum Replacing SGX instructions with IOCTL and system calls, HotCalls 1,200
Total 7,500
@@
Security monitor. We based our implementation on the main branch of SVSM without current privilege level (CPL) support, and added our own support for running user-space enclaves. The security monitor emulates SGX leaf functions based on requests from the NestedSGX-driver. To accommodate these requests, a new type of SVSM protocol called SGX protocol was introduced within the SVSM framework (SVSM, 2022). The security monitor effectively enters a request loop within the SVSM framework, allowing it to continue execution upon the arrival of subsequent requests. The security monitor handles a memory pool acting as the EPC, and faithfully emulates SGX data structures and operations in strict adherence to the SGX specification. We registered customized fault and exception handlers for the enclave, emulating the AEX event. Furthermore, we configured the system call handler by setting the LSTAR MSR to handle the EEXIT leaf function.
@@
With the aforementioned design, the majority of SGX programs can run on NestedSGX without requiring modifications to the source code. All the sample code projects provided by the official Intel SGX SDK, except SampleAEXNotify, ran without any issues on NestedSGX. Furthermore, to simplify the porting of existing applications, we have adapted the Rust SGX SDK (Wang et al., 2019) to the NestedSGX platform. Furthermore, we made slight modifications to the Occlum library OS (Shen et al., 2020) to handle the marshalling buffer and enable its execution on NestedSGX. Building upon Occlum, we implemented the HotCalls optimization (Weisse et al., 2017), allowing certain OCALLs to be processed asynchronously within the enclave without exiting it. [^3] We did not cover all OCALLs but focused on optimizing 15 frequently used ones, such as occlum_ocall_sendmsg, occlum_ocall_recvmsg, occlum_ocall_clock_gettime, occlum_ocall_posix_memalign, occlum_ocall_free, u_sgxprotectedfs_fread_node and u_setsockopt_ocall etc.
Intra-enclave isolation. Since Intel SGX only supports the monolithic model, attempts have been made to support intra-enclave isolation, such as Nested Enclave (Park et al., 2020) and LIGHTENCLAVE (Gu et al., 2022). However, these solutions necessitate hardware modifications that go beyond Intel SGX’s capabilities. We incorporated support for intra-enclave isolation, leveraging the hardware capabilities of memory protection keys for userspace (PKRU). We offer four PKRU APIs as syscalls for enclaves: pkey_set, pkey_alloc, pkey_free, and pkey_mprotect. Additionally, we provide two syscalls that allow enclaves to enable or disable PKRU mechanisms.
Initially, the enclaves invoke the pkru_enable function, informing the security monitor to establish the key manager. Subsequently, the enclaves can use the pkey_alloc and pkey_set functions to obtain a key from the manager and make use of it through pkey_mprotect. On the pkey_mprotect function, the security monitor checks with the key manager to verify if the key has been assigned to the enclaves. If valid, the monitor updates the PKRU-bits of corresponding PTEs by traversing the enclaves’ page tables. The key manager faithfully records and updates key information, including key attributes and the protected virtual address of enclaves. When the enclaves invoke the pkru_free function, the security monitor clears the PKRU-bits of the relevant PTEs. Additionally, the enclaves can call the pkru_disable function to instruct the security monitor to release the key manager.
@@ 5. Security Analysis
While the enclave operates at VMPL0, it is prohibited from directly obtaining the SEV-SNP attestation report. Otherwise, the enclave may incorporate a counterfeit AIK into the report, which could compromise the integrity of the trust chain, particularly if the enclave deliberately exposes the private key of the AIK. This protection is enforced by the fact that SNP_REPORT_REQ can only be initiated from the kernel mode, which is the security monitor. More specifically, SNP_REPORT_REQ necessitates shared memory with the host VMM for passing request and response messages, but we intentionally disable shared memory between the enclave and the host VMM. The hardware also prevents the guest OS (operating at VMPL1) from obtaining an SEV-SNP attestation report for VMPL0. Specifically, the desired VMPL is provided by the guest in the SNP_REPORT_REQ message, and the guest can only generate attestation reports for VMPLs that are greater than or equal to the current VMPL (AMD, 2023d, Sec. 7.3).
@@ 6.1. Experimental setup
We deployed NestedSGX on a server with two AMD EPYC 7543 CPUs (two threads per core, total of 128 logical cores) with 64 GB DDR4 RAM. Every CVM was allocated 4 vCPUs and 4 GB RAM, and ran Ubuntu 22.04 with kernel version 6.5.0-snp-guest (svsm-preview-guest-v3 branch provided by AMD). We configured 512 MB gPA for the security monitor and the enclaves. The modified Linux SGX SDK was based on version 2.20, and the Occlum version was version 0.29.7. We used the main branch of the linux-svsm framework (commitID: 8d518f1). The SVSM framework and the security monitor were compiled with Rust in the default mode (-O0). All programs were compiled with GCC 11.4.0 and the same optimization level (-O2).
The baseline was conducted on the same AMD server running in the SGX SDK simulation mode (sim mode), which does not provide any security guarantees. Specifically, we used the same code base under the same optimization level and compiling options. Notably, as a result of the HotCalls optimization, applications running on Occlum experienced improved performance without the need for any modifications to their source code.
@@ 6.2. Micro-benchmarks
[b]
Table 3. Latencies of emulated SGX leaf instructions. The emulation of ENCLS leaf instructions require VMPL switches, which dominates the cost, while most ENCLU leaf instructions are emulated entirely at VMPL0. The slower performance of ENCLU leaf instructions on NestedSGX is attributed to the use of slower Rust implementations in the cryptographic crates.
[…]
ENCLS ECREATE 3,719 us 8.4 us
EADD 1,421 us 8.0 us
EEXTEND 987 us 33 us
EINIT 811 us 46 us
EREMOVE 1,014 us 7.8 us
EAUG 990 us 7.9 us
EBLOCK 841 us 9.2 us
ELDB/ELDU 1,958 us 9.3 us
EMODPR 1,071 us 9.9 us
EWB 1,819 us 7.9 us
ENCLU EGETKEY 5.0 us 17 us
EREPORT 19 us 30 us
  • $\dagger$
    The vSGX results are taken from the published paper (Zhao et al., 2022).
First, we measured the latencies of emulated SGX leaf instructions and context switches (i.e., ECALLs and OCALLs). Following that, we conducted 4 representative micro-benchmarks: Linux/Unix nbench for memory- and computation-intensive tasks, Lmbench for memory-intensive tasks, WolfSSL for computation-intensive tasks, and FIO for I/O-intensive tasks.
SGX leaf instructions. We measured the emulated SGX leaf instructions in NestedSGX, averaging the results over 10,000 runs. As shown in Table 3, all ENCLS leaf instructions on NestedSGX are significantly faster than vSGX. This is attributed to the requirement of VMPL switches for emulating ENCLS leaf instructions on NestedSGX, whereas vSGX necessitates more costly cross-VM communications. Both EGETKEY and EREPORT on NestedSGX are slower than those on vSGX, which we confirmed is due to the use of the Rust rust-crypto crate that is slower than the implementation used by vSGX. Notably, EGETKEY and EREPORT are not heavily invoked in practice.
[b]
Table 4. Latencies of context switches. The cost of NestedSGX includes the following parts: a VMPL switch round-trip which costs about 19,400 cycles on our platform, the SGX SDK routine which costs 4000–5500 cycles, and the rest is the cost of our unoptimized implementation of the security monitor in emulating SGX instructions (including an EENTER and an EEXIT, which cost about 9,000 cycles). The Intel platform for comparison was equipped with an Intel Xeon E3-1270 v6 CPU and 64 GB RAM.
[…]
ECALL 14,437 cycles (3.8 us) $\approx 1,500$ us 33,584 cycles (12 us)
OCALL 12,586 cycles (3.3 us) - 32,014 cycles (11 us)
  • $\ddagger$
    The vSGX results are taken from the published paper (Zhao et al., 2022).
Context switches. We measured the latency of ECALLs and OCALLs in NestedSGX. The test runs empty edge calls with no explicit parameters 100,000 times and takes the average value. As shown in Table 4, the latency of context switches in NestedSGX is 1.33×1.$3$1.54×1.54$\times$1.54 × higher than Intel SGX, but is still two orders of magnitude faster than vSGX. Notably, existing works on Intel SGX, including EleOS (Orenbach et al., 2017), HotCalls (Weisse et al., 2017) and Switchless calls (Tian et al., 2018) can be further applied to NestedSGX to reduce the cost brought by context switches.
@@
Linux/Unix nbench. Linux/Unix nbench is a benchmark suite that focuses on evaluating the performance of a computer system’s CPU, FPU, and memory system (Mayer, 2017). We utilized an adapted version of nbench, namely SGX-NBench (sgx, 2017) to evaluate the computation performance of NestedSGX. Figure 6(a) demonstrates that NestedSGX introduces an overhead of approximately 1.3% on average over the baseline, which we believe mostly attributes to the context switches introduced by SGX-NBench during the evaluation.
Lmbench. Lmbench is a benchmark suite commonly employed to assess various aspects of system behavior. In our evaluation, we utilized the SGX version of Lmbench, known as SGX-bench (Hasan et al., 2020), specifically employing the bw_mem test to evaluate the memory bandwidth. We conducted 10,000 repetitions of the test and recorded the average values. The memory block size for every access was set to 2 MB. The results, presented in Figure 6(c), indicate that the memory bandwidth achieved is 98.7% of the baseline.
WolfSSL (wol, 2024a). We performed an evaluation of WolfSSL with Intel SGX (wol, 2024b). This benchmark primarily focuses on computation-intensive tasks, such as encryption, decryption, digests, and signature verification. As illustrated in Figure 6(d), NestedSGX introduces an average overhead of about 0.78% compared to the baseline.
Flexible I/O Tester (FIO) (git, 2024) with Occlum. We conducted an evaluation of the I/O performance of NestedSGX by utilizing FIO (v3.28) with Occlum in the default configuration. Our focus was on benchmarking the bandwidth of randomized and sequential read and write operations to the disk, utilizing a single thread. We chose the direct I/O mode for sequential operations. Each test involved accessing a 256 MB file with a block size of 256 KB per access. We repeated the test 10 times, with each iteration lasting 100 seconds, and calculated the average bandwidth from all the runs. As depicted in Figure 6(c), the I/O read performance of NestedSGX is approximately 65.3% of the baseline, while the I/O write performance is similar to the baseline, possibly due to Occlum’s optimizations to reduce the context switches in write operations.